# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix F Using the Performance Difference Lemma
$Q^{*}(s,\pi^{*}(\cdot|s))$ $-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}})\}_{i\in[n]})$
$\leq\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}}),\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\Delta^{i}})$
$\quad\quad-\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}}),\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\Delta^{i}})$
$\quad\quad+\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\pi^{*}(s)_{i},\{\pi^{*}(s)_{j}\}_{j\in\Delta^{k}})$
$\quad\quad-\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\pi^{*}(s)_{i},\{\pi^{*}(s)_{j}\}_{j\in\Delta^{k}})$
$\leq\left|Q^{*}(s,\pi^{*}(s))-\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\pi^{*}(s)_{i},\{\pi^{*}(s)_{j}\}_{j\in\Delta^{i}})\right|$
$\quad\quad+\bigg{|}\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad\quad-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|}$
$\leq\frac{1}{n}\sum_{i\in[n]}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$\quad\quad+\frac{1}{n}\sum_{i\in[n]}\bigg{|}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j,\Delta^{j}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|},$
@@
Lemma F.5.
Fix $s\in\mathcal{S}:=\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$. For each $j\in[n]$, suppose we are given $T$-length sequences of random variables $\{\Delta^{j}_{1},\dots,\Delta^{j}_{T}\}_{j\in[n]}$, distributed uniformly over the support (${[n]\setminus j\choose k-1}$. Further, suppose we are given a fixed sequence $\delta_{1},\dots,\delta_{T}$, where each $\delta_{t}\in(0,1]$ for $t\in[T]$. Then, for each action $a=(a_{g},a_{[n]})=\pi^{*}(s)$, for $t\in[T]$ and $j\in[n]$, define deviation events $B_{t}^{a_{g},a_{j,\Delta_{t}^{j}}}$ such that:
$B_{t}^{a_{g},a_{j,\Delta_{t}^{j}}}\!:=\!\left\{\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{j},a_{g})\!-\!\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{j},a_{g})\right|\!>\!\frac{\sqrt{\frac{n-k+1}{8kn}\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta_{t}}}}{1-\gamma}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}+\epsilon_{k,m}\right\}$ (32)
For $i\in[T]$, we define bad-events $B_{t}$ (which unifies the deviation events) such that
@@
Next, denote $B=\cup_{i=1}^{T}B_{i}$. Then, the probability that no bad event $B_{t}$ occurs is:
@@
$|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})$ $-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})|$
$=\bigg{|}{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})$
$\quad\quad+\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\bigg{|}$
$\leq\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|$
$\quad\quad+\left|\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|$
$\leq\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|+\epsilon_{k,m}$
@@
$\left|\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|$ $\leq\left\|\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right\|_{\infty}$
@@
where the $\epsilon_{k,m}$ follows from Lemma 3.2. From Theorem E.8, we have that with probability at least $1-\delta_{t}$,
$\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|\leq\frac{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta_{t}}}{1-\gamma}\sqrt{\frac{n-k+1}{8kn}}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}$
@@
Here, observe that if we union bound across all the events parameterized by the empirical distributions of $a_{g},a_{j,\Delta_{t}^{j}}\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$ given by $F_{a_{\Delta_{t}^{j}}}$, this also forms a covering of the choice of variables $F_{s_{\Delta_{t}^{j}}}$ by agents $j\in[m]$, and therefore across all choices of $\Delta_{t}^{1},\dots,\Delta_{t}^{n}$ (subject to the permutation invariance of local agents) for a fixed $t$. Thus, from the union bound, we get:
@@
$\Pr[\bar{B}]$ $\geq 1-\sum_{t=1}^{T}\Pr[B_{t}]\geq 1-|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\sum_{t=1}^{T}\delta_{t}$
Lemma F.6.
@@
$\mathbb{E}_{s\sim\mathcal{D}}\left[\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\right]$
$\quad\quad\quad\quad\leq\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
@@
$\mathbb{E}_{s\sim\mathcal{D}}\left[\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\right]$
$=\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\mathbb{E}_{s\sim\mathcal{D}}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
@@
$\quad\mathbbm{1}\left\{\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\leq\frac{\|r_{l}(\cdot,\cdot)\|_{\infty}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}\right\}$
We then study the expected difference between $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))$ and $\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})$.
Observe that:
$\mathbb{E}_{s\sim\mathcal{D}}\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$\quad\quad\quad\quad\quad\quad=\mathbb{E}_{s\sim\mathcal{D}}\left[\mathcal{I}(i,s,k,\delta)\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\right]$
$\quad\quad\quad\quad\quad\quad+\mathbb{E}_{s\sim\mathcal{D}}\left[(1-\mathcal{I}(i,s,k,\delta))\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\right]$

### v3

@@ Appendix G Using the Performance Difference Lemma
$Q^{*}(s,\pi^{*}(\cdot|s))-Q^{*}(s,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}})\}_{i\in[n]})$
$\leq\frac{1}{n}\sum_{i\in[n]}\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}}),\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\Delta^{i}})$
$\quad\quad-\frac{1}{n}\sum_{i\in[n]}\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}}),\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\Delta^{i}})$
$\quad\quad+\frac{1}{n}\sum_{i\in[n]}\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\pi^{*}(s)_{i},\{\pi^{*}(s)_{j}\}_{j\in\Delta^{k}})$
$\quad\quad-\frac{1}{n}\sum_{i\in[n]}\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\pi^{*}(s)_{i},\{\pi^{*}(s)_{j}\}_{j\in\Delta^{k}})$
$\leq\left|Q^{*}(s,\pi^{*}(s))-\frac{1}{n}\sum_{i\in[n]}\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\pi^{*}(s)_{i},\{\pi^{*}(s)_{j}\}_{j\in\Delta^{i}})\right|$
$\quad\quad+\bigg{|}\frac{1}{n}\sum_{i\in[n]}\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad\quad-Q^{*}(s,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|}$
$\leq\frac{1}{n}\sum_{i\in[n]}\left|Q^{*}(s,\pi^{*}(s))-\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$\quad\quad+\frac{1}{n}\sum_{i\in[n]}\bigg{|}\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j,\Delta^{j}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-Q^{*}(s,\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})_{g},\{\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|},$
@@
Lemma G.5.
Fix $s\in\mathcal{S}:=\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$. For each $j\in[n]$, suppose we are given $T$-length sequences of random variables $\{\Delta^{j}_{1},\dots,\Delta^{j}_{T}\}_{j\in[n]}$, distributed uniformly over the support (${[n]\setminus j\choose k-1}$. Further, suppose we are given a fixed sequence $\delta_{1},\dots,\delta_{T}$, where each $\delta_{t}\in(0,1]$ for $t\in[T]$. Let
$\Phi_{k,t}=\frac{1}{1-\gamma}\sqrt{\frac{n-k+1}{8kn}\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta_{t}}}.$
Then, for each action $a=(a_{g},a_{[n]})=\pi^{*}(s)$, for $t\in[T]$ and $j\in[n]$, define deviation events $B_{t}^{a_{g},a_{j,\Delta_{t}^{j}}}$ such that:
$B_{t}^{a_{g},a_{j,\Delta_{t}^{j}}}\!:=\!\left\{\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{j},a_{g})\!-\!\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{j},a_{g})\right|\!>\Phi_{k,t}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}+\epsilon_{k,m}\right\}$ (29)
For $i\in[T]$, we define bad-events $B_{t}$ (which is a union over each deviation event):
@@
Next, denote $B=\bigcup_{i=1}^{T}B_{i}$. Then, the probability that no bad event $B_{t}$ occurs is:
@@
$|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})$ $-\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})|$
$=\bigg{|}{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\hat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})$
$\quad\quad+\hat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\bigg{|}$
$\leq\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\hat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|$
$\quad\quad+\left|\hat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|$
$\leq\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\hat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|+\epsilon_{k,m}$
@@
$\left|\hat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|$
$\leq\left\|\hat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right\|_{\infty}$
@@
where the $\epsilon_{k,m}$ follows from Lemma 4.3. From Theorem F.8, we have that with probability at least $1-\delta_{t}$,
$\bigg{|}{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})$ $-\hat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\bigg{|}\leq\Phi_{k,t}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}$
$=\frac{1}{1-\gamma}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta_{t}}}\sqrt{\frac{n-k+1}{8kn}}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}$
@@
Here, observe that if we union bound across all the events parameterized by the empirical distributions of $a_{g},a_{j,\Delta_{t}^{j}}\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$ given by $F_{a_{\Delta_{t}^{j}}}$, this also forms a covering of the choice of variables $F_{s_{\Delta_{t}^{j}}}$ by agents $j\in[m]$, and therefore across all choices of $\Delta_{t}^{1},\dots,\Delta_{t}^{n}$ (subject to the permutation invariance of local agents) for a fixed $t$.
Thus, from the union bound, we get:
@@
$\Pr[\bar{B}]$ $\geq 1-\sum_{t=1}^{T}\Pr[B_{t}]\geq 1-|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\sum_{t=1}^{T}\delta_{t},$
which proves the claim.∎
Lemma G.6.
@@
$\mathbb{E}_{s\sim\mathcal{D}}\bigg{[}\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\bigg{|}Q^{*}(s,\pi^{*}(s))$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}\bigg{]}$
$\quad\quad\quad\quad\leq\frac{\tilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\tilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
@@
$\mathbb{E}_{s\sim\mathcal{D}}\bigg{[}\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\bigg{|}Q^{*}(s,\pi^{*}(s))$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}\bigg{]}$
$=\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\mathbb{E}_{s\sim\mathcal{D}}\bigg{|}Q^{*}(s,\pi^{*}(s))$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\hat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
Let
$\Phi_{k,\delta}=\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}.$
@@
$\mathbbm{1}\left\{\left|Q^{*}(s,\pi^{*}(s))-\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\leq\frac{\|r_{l}(\cdot,\cdot)\|_{\infty}}{1-\gamma}\Phi_{k,\delta}+\epsilon_{k,m}\right\}$
The expected difference between $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))$ and $\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})$ is bounded as follows:
$\mathbb{E}_{s\sim\mathcal{D}}\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$=\mathbb{E}_{s\sim\mathcal{D}}\left[\mathcal{I}(i,s,k,\delta)\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\right]$
$+\mathbb{E}_{s\sim\mathcal{D}}\left[(1-\mathcal{I}(i,s,k,\delta))\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\right]$